OpenStudy (anonymous):

–6p + 7 = 3(2p – 3) – 4(–10 + 4p)

OpenStudy (anonymous):

-6p + 7 = 3(2p - 3) - 4(-10 + 4p) -6p + 7 = 6p - 9 + 40 - 16p -6p + 7 = -10p + 31 4p + 7 + 31 4p = 24 p = 6

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

no problem